OpenStudy (anonymous):

Find the absolute maximum and absolute minimum values of f on the given interval. f(x)=ln(x^2+x+1), [-1,1]

OpenStudy (campbell_st):

well you can differentiate and find the stationary point(s)... I think the stationary point is a min... and text the endpoints for max values...

OpenStudy (anonymous):

idk how to work it out completely

OpenStudy (campbell_st):

ok... so the end point substitute x = -1 and x = 1 into the equation... then the derivative is \[\frac{dy}{dx} = \frac{2x + 1}{x^2 + x + 1}\] so the stationary point occurs when 2x + 1 = 0 hope this helps

OpenStudy (anonymous):

The answer in the back of the book says f(1)=ln3, f(-1/2)= ln 3/4... don't understand

OpenStudy (campbell_st):

so a stationary point occurs at x = -1/2 test at x = -1/4 and x = -3/4 for a change in the slope of the tangent... x = -1/4 dy/dx > 0 x = -3/4 dy/dx < 0 so x = -1/2 is a minimum substitute it to find the value hope this makes sense

OpenStudy (campbell_st):

f(1) = ln(3) is the max value of the curve between [-1, 1] f(-1/2) = ln(3/4) is the minimum value of the curve in the same domain [-1,1]

OpenStudy (campbell_st):

the cuve looks like